# Locating and Editing Factual Associations in Mamba

### v1

@@ 5 Linearity of Relation Embedding (LRE)
With activation patching we can identify where facts are located inside a LM. We are also interested in understanding how LM extracts this information given $x=(s,r)$. Figures 2(b) and 5 show a clear separation of roles in early-mid and later layers in Mamba. We observe a similar phenomena in autoregressive transformer LMs (Meng et al., 2022a; b; Geva et al., 2023). According to Geva et al. (2023), in transformer LMs, the subject entity representation $\mathbf{s}$, at the subject last token position, goes through an enrichment process, mediated by the MLP in the early-mid layers, where $\mathbf{s}$ is populated with different facts/attributes relevant to the subject entity $s$. And, then at the last token position attention modules perform a query in the enriched $\mathbf{s}$ to extract the answer to the prompt $x=(s,r)$. Hernandez et al. (2023) approximate the query performed in the enriched $\mathbf{s}$ for a specific relation $r$ by taking the first order Taylor series approximation (Lre) of the LM computation $F$ as -
@@
Hernandez et al. (2023) show that for a range of different relations it is possible to achieve a Lre that is faithful to the model computation $F$ by averaging the approximations of J and $b$ calculated on just $n=5$ examples. We utilize Lre to understand the complexity of decoding factual relations in Mamba. We find the hyperparameters $\beta$, $\rho$ and the layer $\ell$ (where to extract enriched $\mathbf{s}$ from) using grid search. For mathematical and implementation details, see Hernandez et al. (2023).
We plot the faithfulness of Lre with $n=5$ samples on Figure 7. faithfulness represents the portion of facts in $r$ that can be correctly retrieved if the LM computation $F(\mathbf{s},r)$ is replaced with $\textsc{Lre}(\mathbf{s})$.
@@
We only calculate Lre for the factual relations in the Relations dataset. Figure 7 shows that only for $10$ out of $26$ factual relations a linear Lre can achieve more than $50\%$ faithfulness. For comparison, in the same sized Pythia-2.8b Lre achives $>50\%$ faithfulness for $11$ factual relations (see Appendix E). And, in both Mamba and Pythia, Lre fails to achieve good faithfulness for the relations where the range (the number of unique answers) is larger.
@@ 6 Attention Knock-out in Mamba?
In transformer LMs, attention modules bring information from the other token positions. Each of the attention heads in an attention module $attn^{(\ell)}$ calculates an attention matrix L, where $\text{L}_{k,q}$ quantifies how much attention is being paid to the $q^{th}$ token by the $k^{th}$ token. See Vaswani et al. (2017) for details on the attention operation. We can block information flow from $q^{th}$ token to $k^{th}$ via a specific attention head by simply setting $\text{L}_{k,q}:=-\infty$ in the forward pass. This patch-dependent attention blocking (or attention “knock-out”) is also a form of causal mediation analysis and it has been effective in understanding the information flow in transformer LMs (Geva et al., 2023; Wang et al., 2022; Todd et al., 2023).
For Mamba, Ali et al. (2024) show that the amount of information retained in the $k^{th}$ token state $s_{k}^{(\ell)}$, from the convolved state at $q^{th}$ token $c_{q}^{(\ell)}$ (where $q<k$), after the selective-SSM operation (see Equations 4 and 5) can be visualized as an attention matrix per channel (dim). Since the selective-SSM operation is linear, the information retained in $s_{k}^{(\ell)}$ from $c_{q}^{(\ell)}$ can be calculated accurately as $\tilde{\alpha}_{k,q}=\overline{\text{C}}_{k}^{(\ell)}\Big{(}\prod_{i=q+1}^{k}\overline{\text{A}}_{i}^{(\ell)}\Big{)}\overline{\text{B}}_{q}^{(\ell)}c_{q}^{(\ell)}$, where $\overline{\text{A}}_{i}^{(\ell)}$, $\overline{\text{B}}_{i}^{(\ell)}$, and $\overline{\text{C}}_{i}^{(\ell)}$ are input-dependent parameters for the $i^{th}$ token. See Gu & Dao (2023) and Ali et al. (2024) for details. We ask: can we block the information flow from the $q^{th}$ token to the $k^{th}$ token in Mamba by subtracting out $\tilde{\alpha}_{k,q}$ from $s_{k}^{(\ell)}$? If so, attention knockout experiments in Mamba become feasible.
We find that blocking information flow from the $q^{th}$ token to the $k^{th}$ token in the Conv + SSM operation can be difficult. Note that, since $c_{q}^{(\ell)}$ is a convolved state with a receptive field of size 4 in Mamba-2.8b, the states $c_{q+1}^{(\ell)}$, $c_{q+2}^{(\ell)}$, and $c_{q+3}^{(\ell)}$ retain information from $a_{q}^{(\ell)}$. And, these states can “leak” information about $a_{q}^{(\ell)}$ to $s_{k}^{(\ell)}$. To stop this leakage, we would want to subtract from $s_{k}^{(\ell)}$ all the information retained about $a_{q}^{(\ell)}$. However, calculating this is challenging because of the SiLU non-linearity after Conv1D (see Equation 4).
@@
In Appendix D, we present a version of attention knock-out experiment, where we block information flow of the $q^{th}$ token to all the future tokens with Conv + SSM operation. We recognize that this intervention is not as surgical as blocking information from $q^{th}$ token to just $k^{th}$ token. However, with some caveats, this experiment leads us to draw some conclusion about the information flow in Mamba that is similar to what Geva et al. (2023) found for GPT LMs. See Appendix D for details.
@@ 7 Related Works
Locating Factual Knowledge in Language Models.
To make factually correct statements about the world a LM has to store factual knowledge about real world entities somewhere in its parameters. Understanding how and where a neural network stores knowledge is a core problem for interpretability and it has thus been studied from several perspectives (Ji et al., 2021; Wang et al., 2014). One line of work trains classifiers to probe for properties encoded in model representations (Ettinger et al., 2016; Shi et al., 2016; Hupkes et al., 2018; Conneau et al., 2018; Belinkov et al., 2017; Belinkov & Glass, 2019). However, the flexibility of these classifiers can lead to overestimating model knowledge and capabilites (Belinkov, 2022). Causal mediation analysis methods (Pearl, 2022) attempt to measure the causal contribution of intermediate states to task performance. Meng et al. (2022a; b) use activation patching to identify key MLP modules for factual recall, highlighting the middle layers at particular token positions as being especially important. Similarly, Geva et al. (2023) apply causal mediation analysis to attention modules, seeking to understand the mechanism of cross-token factual information flow inside transformer LMs.
@@ 8 Discussion
In this paper we have explored whether the analytical methods and tools developed for transformer LMs can also be applied on Mamba, a recurrent state-space architecture, and we have compared the mechanisms of factual recall between the two classes of models. Our experiments have been limited to Mamba-2.8b, the largest available LM of that family, and comparisons to Pythia-2.8b (selected over Llama or GPT models as they do not offer pretrained models of similar size).
With activation patching we have found that, similar to autoregressive transformer LMs, Mamba shows signs of localization at the last subject token and at specific layer ranges while recalling a fact. We have applied rank-one model editing (ROME) to rewrite a fact in specific parameters in Mamba. Although, unlike transformers, Mamba has no MLP modules, we find that their $\text{W}_{o}$ weights can receive edits with good generalization and specificity at a range of layers. Then we have studied the linearity of the embeddings of factual relations in Mamba and have found that many can be well approximated by Lre, again resembling transformer LMs. Although Mamba has no attention modules, we have been able to partially adapt the tools of attention knock-out using an SSM blocking technique, and this reveals information flows similar to transformers. In summary, we find that many of the tools used to interpret and edit large transformers can be adapted to work with Mamba as well, and that despite the architectural differences, when it comes to factual recall, Mamba shares many similarities with transformer LMs.
@@
We ran all experiments on workstations with either 80GB NVIDIA A100 GPUs or 48GB A6000 GPUs, using the HuggingFace Transformers library (Wolf et al., 2019) and PyTorch (Paszke et al., 2019). We make use of publicly available datasets CounterFact and Relations in this work. We will make our source code open source after the anonymity period has ended.
@@ A.1 CounterFact
Meng et al. (2022a) developed the CounterFact dataset for evaluating the efficacy of counterfactual edits in language models. It was prepared by adapting ParaRel (Elazar et al. (2021)) and scraping Wikidata. The dataset contains $21,919$ requests $\{s,r,o,o^{*},\pi^{*}\}$ where $o$ is the correct answer to the prompt $x=(s,r)$, $o^{*}$ is the counterfactual edit request, and $\pi^{*}\sim\mathcal{P}(s,r)$ is a paraphrase of the prompt $x=(s,r)$ to test for generalizability (PS). Each of the records also contain some neighborhood prompts $\pi_{N}$ to test for specificity (NS) and some generation prompts $\pi_{G}$ to test if LM generation post-edit is fluent and consistent with the edit. Please refer to Meng et al. (2022a) for details on the curation of this dataset.
We evaluate ROME performance in Mamba-2.8b and Pythia-2.8b on the first $2000$ records from CounterFact.
@@ A.2 Relations
The Relations dataset introduced in Hernandez et al. (2023) consists of $47$ relations of $4$ types; factual, linguistic, bias, and commonsense. A relation $r$ is an association between two entities. For example, the relation, $r=\textit{professionally played the sport}$ connects the subject $s=\textit{Michael Jordan}$ with the object $o=\textit{basketball}$. The dataset contains a set of $(s,o)$ for each relation $r$.
@@
In the scope of this paper, we only utilize the $26$ factual relations from this dataset. We evaluate Lre in Mamba and Pythia for all the $26$ factual relations. We also use this dataset for locating key fact-mediating states in Section 3 and Appendix C. We randomly sample 400 examples $(s,r,o)$ across 6 different factual relations - place in city, country capital city, person occupation, plays pro sport, company hq, and product by company. For each of these examples we randomly select another example within the same relation $(s^{*},r,o^{*})$ such that $s\neq s^{*}$ and $o\neq o^{*}$. The average indirect effect (IE) of applying activation patching over these 400 examples is depicted on Figures 2(b), 3, 5 for Mamba-2.8b) and on Figure 9 (for Pythia-2.8b).
Appendix B Isolating The Contribution of $\text{W}_{o}^{(\ell)}$
@@
Recall from Figure 1 and Equation 2 that when $o_{i}^{(\ell)}$ is restored, the $s_{i}^{(\ell)}$ and $g_{i}^{(\ell)}$ are restored as well. To isolate the contribution of only $\text{W}_{o}^{(\ell)}$ we subtract out $\text{IE}_{s_{i}^{(\ell)}}+\text{IE}_{g_{i}^{(\ell)}}$ from $\text{IE}_{o_{i}^{(\ell)}}$ and plot the results on Figure 8. Notice that, subtracting $\text{IE}_{s_{i}^{(\ell)}}$ cancels out the high indirect effect at the late site shown by later layers at the last token position. But, together $\text{IE}_{s_{i}^{(\ell)}}+\text{IE}_{g_{i}^{(\ell)}}$ cannot cancel out high $\text{IE}_{o_{i}^{(\ell)}}$ of early-mid layers at the last subject token, reconfirming the mediating role of $\text{W}_{o}^{(\ell)}$ in those layers while recalling a fact.
@@
Figure 8: Isolating the contribution of $\text{W}_{o}^{(\ell)}$. $\text{IE}_{s_{i}^{(\ell)}}+\text{IE}_{g_{i}^{(\ell)}}$ subtracted from $\text{IE}_{o_{i}^{(\ell)}}$. Notice that $\text{IE}_{o_{i}^{(\ell)}}-\big{(}\text{IE}_{s_{i}^{(\ell)}}+\text{IE}_{g_{i}^{(\ell)}}\big{)}$ still shows higher causal effect at the earlier site (more pronounced than $\text{IE}_{g_{i}^{(\ell)}}$) while the high causal effect at the later site cancels out.
Appendix C Locating Key Modules in Pythia-2.8b
@@
Figure 9: Average indirect effect of residual state, MLP, and attention outputs in Pythia-2.8b over 400 facts. For MLP and attention outputs a window of 10 layers around $\ell$ is restored as restoring for just one layer barely show visible patterns.
@@
Appendix D Blocking Information Propagation With $s_{i}$ Paths
In Section 6, we discuss why surgically performing attention knock-out experiments on Mamba is challenging due to certain architectural choices - we cannot block information flow from the $q^{th}$ token to only the $k^{th}$ token because of the SiLU non-linearity in Equation 4. However, for a layer $\ell$ we can block information propagation via Conv + SSM from the $q^{th}$ token to all the future tokens ($k>q$) by mean ablating $a_{q}^{(\ell)}$. For a window of 10 layers around $\ell$ we block-out information propagation of the subject, non-subject, and the prompt-last token positions. The results plotted in Figure 11 show relative change in $p(o)$ by blocking out $q^{th}$ token across different layers as $\nicefrac{{\Big{(}p\big{(}o\,|\,a_{q}^{(\ell)}:=\mathbb{E}\big{[}a^{(\ell)}\big{]}\big{)}-p(o)\Big{)}}}{{p(o)}}$.
Figure 11: Relative change in $p(o)$ when information flow to future tokens via $s_{i}$ paths is blocked from either subject, non-subject, or the prompt_last token positions. For each layer $\ell$, $s_{i}$ paths were blocked for a window of 10 layers around $\ell$.
Figure 11 leads us to draw similar conclusions about information flow in Mamba as Geva et al. (2023) observe in transformer LMs.
  • (a)
    The purple lines show that blocking out non-subject information flow in early middle layers can bring down $p(o)$ by up to $50\%$. Non-subject tokens are used to specify the relation $r$. This observation leads us to believe that Mamba propagates relation specific information to future tokens using Conv+SSM operation in early-middle layers
  • (b)
    Interestingly, the green lines (blocking the subject information flow) shows two valleys -
    1. 1.
    2. 2.
With this experiment, we cannot make strong claims on the role of the last token position (prompt-last) like Geva et al. (2023). As we block out information to all future tokens, the intermediate states in between the ablated token and the last token are getting corrupted as well.
@@ Appendix D Blocking Information Propagation With $s_{i}$ Paths
Appendix F Lre Performance Over Different Relations

### v2

@@ 5 Linearity of Relation Embedding (LRE)
With activation patching we can identify where facts are located inside a LM. We are also interested in understanding how LMs extract this information given $x=(s,r)$. Figures 2(b) and 5 show a clear separation of roles in early-mid and later layers in Mamba. We observe a similar phenomenon in autoregressive transformer LMs (Meng et al., 2022a; b; Geva et al., 2023). According to Geva et al. (2023), in transformer LMs, the subject entity representation $\mathbf{s}$, at the subject last token position, goes through an enrichment process, mediated by the MLP in the early-mid layers, where $\mathbf{s}$ is populated with different facts/attributes relevant to the subject entity $s$. Then, at the last token position, attention modules perform a query on the enriched $\mathbf{s}$ to extract the answer to the prompt $x=(s,r)$. Hernandez et al. (2023) approximate the query operation performed on the enriched $\mathbf{s}$ for a specific relation $r$ by taking the first order Taylor series approximation (Lre) of the LM computation $F$ as
@@
Hernandez et al. (2023) show that for a range of different relations it is possible to achieve a Lre that is faithful to the model computation $F$ by averaging the approximations of J and $b$ calculated on just $n=5$ examples. We utilize Lre to understand the complexity of decoding factual relations in Mamba. We find the hyperparameters $\beta$, $\rho$ and the layer $\ell$ (where to extract the enriched $\mathbf{s}$ from) using grid search. For mathematical and implementation details, see Hernandez et al. (2023).
We plot the faithfulness of Lre with $n=5$ samples on Figure 7. The metric faithfulness represents the portion of facts $(s,r,o)$ that can be correctly retrieved if the LM computation $F(\mathbf{s},r)$ is replaced with $\textsc{Lre}(\mathbf{s})$, a simple affine transformation.
@@
We only calculate Lre for the factual relations in the Relations dataset. Figure 7 shows that only for $10$ out of $26$ factual relations can a linear Lre achieve more than $50\%$ faithfulness. For comparison, in the same sized Pythia-2.8b Lre achives $>50\%$ faithfulness for $11$ factual relations (see Appendix E). And, in both Mamba and Pythia, Lre fails to achieve good faithfulness for the relations where the range (the number of unique answers) is large. These findings align with what Hernandez et al. (2023) observed on GPT and LLaMA models; suggesting that, similar to transformer LMs, factual knowledge might be heterogeneously represented for different relations in Mamba.
@@ 6 Attention Knock-out in Mamba?
Attention modules mediate the flow of information across different token positions in transformer LMs. In attention “knock-out” experiments the information that flows through a specific edge (from $k^{th}$ token to $q^{th}$ token) via a certain attention head is blocked to understand if critical information flows through that edge. This is also a form of causal mediation analysis and it has been effective in understanding the information flow in transformer LMs (Geva et al., 2023; Wang et al., 2022; Todd et al., 2023). In Mamba, information from past tokens is retained in the $s_{i}$ states, with the Conv + SSM operations (see Figure 1 and Equations 35). We ask, can we perform experiments similar to attention knock-out experiments in Mamba in order to understand how it moves factual information?
@@
We find that performing similar experiments in Mamba can be difficult. The use of Conv with a non-linearity in conjunction with selective-SSM make it challenging to remove the information retained in the $q^{th}$ token from the $k^{th}$ token (see Appendix B for details). However, it is possible to block the propagation of information from the $k^{th}$ token to all the future tokens via Conv + SSM operation by mean-ablation. Specifically, for a layer $\ell$, we set $a_{k}^{(\ell)}:=\mathbb{E}\big{[}a^{(\ell)}\big{]}$, where $\mathbb{E}\big{[}a^{(\ell)}\big{]}$ is the mean of $a^{(\ell)}$ states collected with 10,000 tokens from WikiText-103 by Merity et al. (2016). We recognize that this intervention may not be as surgical as cutting a specific edge. However, with some caveats, this experiment suggests that the factual information flow in Mamba is similar to what Geva et al. (2023) observed in GPT LMs.
We randomly sample 700 facts across 6 factual relations from the Relations dataset. For each of those examples we block-out information propagation of the subject, non-subject, and the prompt-last token positions for a window of $10$ layers around a specific layer $\ell$. The effect of blocking out Conv + SSM information flow for certain layer-token ($\ell-k$) positions is measured as the relative change in $p(o)$ with $\nicefrac{{\Big{(}p\big{(}o\,|\,a_{k}^{(\ell)}:=\mathbb{E}\big{[}a^{(\ell)}\big{]}\big{)}-p(o)\Big{)}}}{{p(o)}}$. Figure 8 shows the averaged result and it leads us to draw the following conclusions about how factual information flows in Mamba:
Figure 8: Relative change in $p(o)$ when information flow from $a_{k}^{(\ell)}$ to future tokens via $s_{i}$ paths is blocked, with $k$ taking the value of either subject, non-subject, or the prompt_last token positions. For each layer $\ell$, $s_{i}$ paths were blocked for a window of 10 layers around $\ell$.
  • (a)
    The purple lines show that blocking out non-subject information flow in early middle layers can bring down $p(o)$ by up to $50\%$. Non-subject tokens are used to specify the relation $r$. This observation leads us to believe that Mamba propagates relation specific information to future tokens using Conv+SSM operations in early-middle layers.
  • (b)
    Interestingly, the green lines (blocking the subject information flow) shows two valleys:
    1. 1.
    2. 2.
  • (c)
    The blue dashed lines indicate the effect of blocking the information of only the subject last token. If the ablation is performed in very early layers, later layers can start to compensate for that. However, the valley around layers 20-21 suggests that Mamba expects to recognize the full subject entity by then in order to recall relevant associations (enrichment). Notably, activation patching results for $o_{i}$ and $s_{i}$—states that we hypothesize take crucial part in the enrichment process—also show strong indirect effect around that region (Figures 3a, 3b, and 5a). The blue line follows the green line after layer 30. The weaker effect observed might be because ablating subject last token is not always enough to remove all the subject information. For example, in Eiffel Tower, Eiffel (tokenized as E, iff, el) is more informative than the last token Tower.
These findings align with how factual information flows through attention modules in autoregressive transformer LMs, as observed by Geva et al. (2023) in GPT. However, unlike Geva et al. (2023), we cannot make strong claims about the unique role of the final token position (prompt-last) with this experiment. As we block out information flow to all future tokens, the intermediate states in between the ablated $k^{th}$ token and the last token are affected as well.
@@ 7 Related Works
Locating Factual Knowledge in Language Models. To make factually correct statements about the world, a LM has to store factual knowledge about real world entities somewhere in its parameters. Understanding how and where a neural network stores knowledge is a core problem for interpretability and it has thus been studied from several perspectives (Ji et al., 2021; Wang et al., 2014). One line of work trains classifiers to probe for properties encoded in model representations (Ettinger et al., 2016; Shi et al., 2016; Hupkes et al., 2018; Conneau et al., 2018; Belinkov et al., 2017; Belinkov & Glass, 2019). However, the flexibility of these classifiers can lead to overestimating model knowledge and capabilites (Belinkov, 2022). Causal mediation analysis methods (Pearl, 2022) attempt to measure the causal contribution of intermediate states to task performance. Meng et al. (2022a; b) use activation patching to identify key MLP modules for factual recall, highlighting the middle layers at particular token positions as being especially important. Similarly, Geva et al. (2023) apply causal mediation analysis to attention modules, seeking to understand the mechanism of cross-token factual information flow inside transformer LMs.
@@ 8 Discussion
In this paper we have set out to understand whether the analytical methods and tools developed for transformer LMs can also be applied on the Mamba recurrent state-space architecture. Although our experiments have been limited to Mamba-2.8b, the largest available LM of that family, and comparisons to the similarly-sized transformer Pythia-2.8b, the methods we have introduced are general, and can be used to analyze to any state-space model.
[…]
The similarity that we have observed between factual recall mechanisms in transformers and Mamba leads us to speculate that the autoregressive language modeling task itself induces a pattern of localized factual recall that is independent of modeling architecture. When constraining a model to process text from beginning to end, the ordering creates a specific bottleneck in the information flows: the end of a subject becomes a singular moment at which recognition of the subject is both possible and useful, and we find that both transformers and Mamba arrange their computations to localize factual recall at that moment. We hypothesize that other future autoregressive LMs architectures should expect to see similar locality in factual recall as well.
In summary, we find that many of the tools used to interpret and edit large transformers can be adapted to work with Mamba, and we are optimistic that those tools will continue to be useful as architectures continute to evolve.
@@
We ran all experiments on workstations with either 80GB NVIDIA A100 GPUs or 48GB A6000 GPUs, using the HuggingFace Transformers library (Wolf et al., 2019) and PyTorch (Paszke et al., 2019). We make use of publicly available datasets CounterFact and Relations in this work.
Acknowledgements
This research has been supported by a grant from Open Philanthropy (DB, AS), and an NSF Computer and Information Science and Engineering Graduate Fellowship Fellowship (DA). We are also grateful to the Center for AI Safety (CAIS) for sharing their compute resources, which supported many of our experiments. Some of our initial analyses were conducted with a beta version of NNsight (Fiotto-Kaufman et al., 2024) on an implementation of Mamba instrumented for research by Jaden Fiotto-Kaufmann.
@@ A.1 CounterFact
Meng et al. (2022a) developed the CounterFact dataset for evaluating the efficacy of counterfactual edits in language models. It was prepared by adapting ParaRel (Elazar et al. (2021)) and scraping Wikidata333www.wikidata.org/wiki/Wikidata:Main_Page. The dataset contains $21,919$ requests $\{s,r,o,o^{*},\pi^{*}\}$ where $o$ is the correct answer to the prompt $x=(s,r)$, $o^{*}$ is the counterfactual edit request, and $\pi^{*}\sim\mathcal{P}(s,r)$ is a paraphrase of the prompt $x=(s,r)$ to test for generalizability (PS). Each of the records also contain some neighborhood prompts $\pi_{N}$ to test for specificity (NS) and some generation prompts $\pi_{G}$ to test if LM generation post-edit is fluent and consistent with the edit. Please refer to Meng et al. (2022a) for details on the curation of this dataset.
We evaluate ROME performance in Mamba-2.8b (Figure 6a) and Pythia-2.8b (Figure 6b) on the first $2000$ records from CounterFact.
@@ A.2 Relations
The Relations dataset introduced in Hernandez et al. (2023) consists of $47$ relations of $4$ types: factual, linguistic, bias, and commonsense. A relation $r$ is an association between two entities. For example, the relation, $r=\textit{professionally played the sport}$ connects the subject $s=\textit{Michael Jordan}$ with the object $o=\textit{basketball}$. The dataset contains a set of $(s,o)$ for each relation $r$.
@@
In the scope of this paper, we only utilize the $26$ factual relations from this dataset. We evaluate Lre in Mamba and Pythia for all the $26$ factual relations. We also use this dataset for locating key fact-mediating states in Section 3 and Appendix D. We randomly sample 400 examples $(s,r,o)$ across 6 different factual relations - place in city, country capital city, person occupation, plays pro sport, company hq, and product by company. For each of these examples we randomly select another example within the same relation $(s^{*},r,o^{*})$ such that $s\neq s^{*}$ and $o\neq o^{*}$. The average indirect effect (IE) of applying activation patching over these 400 examples is depicted on Figures 2(b), 3, 5 for Mamba-2.8b) and on Figure 10 (for Pythia-2.8b). We use the same set of $6$ relations in Section 6 where we adapt attention knock-out to Mamba.
Appendix B Challenges in Performing Attention Knock-out in Mamba
Attention heads in autoregressive transformer LMs and Conv + SSM operations in Mamba play a similar role: bringing/retaining information from the past tokens. Attention “knock-out” is a type of causal mediation analysis that tries to understand information flow in transformer LMs by cutting off information propagation from $k^{th}$ token to $q^{th}$ token position. In transformers, each of the attention heads in an attention module $attn^{(\ell)}$ calculates an attention matrix L, where $\text{L}_{q,k}$ quantifies how much attention is being paid to the $k^{th}$ token by the $q^{th}$ token with this specific attention head (see Vaswani et al. (2017) for details on the attention operation). We can block the information flow from $k^{th}$ token to $q^{th}$ token via a specific attention head by simply setting $\text{L}_{q,k}:=-\infty$ in the forward pass.
For Mamba, Ali et al. (2024) show that the amount of information retained in the $q^{th}$ token state $s_{q}^{(\ell)}$, from the convolved state at $k^{th}$ token $c_{k}^{(\ell)}$ (where $k<q$), after the selective-SSM operation (see Equations 4 and 5) can be visualized as an attention matrix per channel. Since the selective-SSM operation is linear, the information retained in $s_{q}^{(\ell)}$ from $c_{k}^{(\ell)}$ can be calculated accurately as $\tilde{\alpha}_{q,k}^{(\ell)}=\overline{\text{C}}_{q}^{(\ell)}\Big{(}\prod_{i=k+1}^{q}\overline{\text{A}}_{i}^{(\ell)}\Big{)}\overline{\text{B}}_{k}^{(\ell)}c_{k}^{(\ell)}$, where $\overline{\text{A}}_{i}^{(\ell)}$, $\overline{\text{B}}_{i}^{(\ell)}$, and $\overline{\text{C}}_{i}^{(\ell)}$ are input-dependent parameters for the $i^{th}$ token. See Gu & Dao (2023) and Ali et al. (2024) for details on selective-SSM operation. We ask: can we block the information flow from the $k^{th}$ token to the $q^{th}$ token in Mamba by subtracting out $\tilde{\alpha}_{q,k}^{(\ell)}$ from $s_{q}^{(\ell)}$? If so, attention knockout experiments in Mamba become feasible.
We find that blocking information flow via Conv + SSM operation through this specific edge from the $k^{th}$ token to the $q^{th}$ token can be challenging in Mamba. Note that, since $c_{k}^{(\ell)}$ is a convolved state with a receptive field of size 4 in Mamba-2.8b, the states $c_{k+1}^{(\ell)}$, $c_{k+2}^{(\ell)}$, and $c_{k+3}^{(\ell)}$ also retain information from $a_{k}^{(\ell)}$. Which means that even if we subtract $\tilde{\alpha}_{q,k}^{(\ell)}$ from $s_{q}^{(\ell)}$, these states can “leak” information about $a_{k}^{(\ell)}$ to $s_{q}^{(\ell)}$. To stop this leakage, we would want to subtract from $s_{q}^{(\ell)}$ all the information retained from $a_{k}^{(\ell)}$ via $c_{k+1}^{(\ell)}$, $c_{k+2}^{(\ell)}$, and $c_{k+3}^{(\ell)}$ states as well. However, accurately calculating this is challenging because of the SiLU non-linearity after Conv1D (see Equation 4).
In our initial experiments we tested subtracting only $\tilde{\alpha}_{q,k}^{(\ell)}$ from $s_{q}^{(\ell)}$. But we found that Mamba-2.8b could often refer to the $k^{th}$ token from the $q^{th}$ token in copy and factual recall tasks.
Appendix C Isolating The Contribution of $\text{W}_{o}^{(\ell)}$
@@
Recall from Figure 1 and Equation 2 that when $o_{i}^{(\ell)}$ is restored, the $s_{i}^{(\ell)}$ and $g_{i}^{(\ell)}$ are restored as well. To isolate the contribution of only $\text{W}_{o}^{(\ell)}$ we subtract out $\text{IE}_{s_{i}^{(\ell)}}+\text{IE}_{g_{i}^{(\ell)}}$ from $\text{IE}_{o_{i}^{(\ell)}}$ and plot the results on Figure 9. Notice that subtracting $\text{IE}_{s_{i}^{(\ell)}}$ cancels out the high indirect effect at the late site shown by later layers at the last token position. But, together $\text{IE}_{s_{i}^{(\ell)}}+\text{IE}_{g_{i}^{(\ell)}}$ cannot cancel out high $\text{IE}_{o_{i}^{(\ell)}}$ observed at the early site, that is early-mid layers at the last subject token. This reconfirms the mediating role of $\text{W}_{o}^{(\ell)}$ at the early site while recalling a fact.
@@
Figure 9: Isolating the contribution of $\text{W}_{o}^{(\ell)}$. $\text{IE}_{s_{i}^{(\ell)}}+\text{IE}_{g_{i}^{(\ell)}}$ subtracted from $\text{IE}_{o_{i}^{(\ell)}}$. Notice that $\text{IE}_{o_{i}^{(\ell)}}-\big{(}\text{IE}_{s_{i}^{(\ell)}}+\text{IE}_{g_{i}^{(\ell)}}\big{)}$ still shows higher causal effect at the early site (more pronounced than $\text{IE}_{g_{i}^{(\ell)}}$) while the high causal effect at the late site cancels out.
Appendix D Locating Key Modules in Pythia-2.8b
@@
Figure 10: Average indirect effect of residual state, MLP, and attention outputs in Pythia-2.8b over 400 facts. For MLP and attention outputs a window of 10 layers around $\ell$ is restored, as restoring just one layer barely shows visible patterns.
@@
Appendix F Lre Performance Across Different Relations